OpenStudy (anonymous):

log(low "3")((x^2)+3)=4

OpenStudy (anonymous):

asking for the solution set

OpenStudy (anonymous):

is it log(log 3) ???

OpenStudy (anonymous):

no just log with a 3 to the right but its lower than the log

OpenStudy (anonymous):

i have got the answer...bear with me...

OpenStudy (anonymous):

log(low "3")((x^2)+3)=4 ln(x^2 +3) / ln 3 = 4 ln(x^2 +3) = 4 * ln 3 ln(x^2 +3) = ln 3^4= ln 81 x^2+3 = 81 x^2= 78 x=+- sqrt (78)

OpenStudy (anonymous):

i also got same answer

OpenStudy (anonymous):

was that correct....i m not sure? :(

OpenStudy (anonymous):

ah ...someone is with me !!!

OpenStudy (anonymous):

should be correct then

OpenStudy (anonymous):

yah right. I was actually solving it and happy to see same answer with you

OpenStudy (anonymous):

gud stuff